the length of a rectangle is 3in. less than twice its width. If the perimeter of the rectangle is 145 in. find the dimension of the rectangle.
Answer by checkley71(8403)

L=2W-3
2W+2(2W-3)=145
2W+4W-6=145
6W=145+6
6W=151
W=151/6
W=25.167 IN. IS THE WIDTH.
2*25.167-3=47.33 IN. IS THE LENGTH.
PROOF:
2*25.167+2*47.33=145
50.33+94.67=145
145=145